Write Mayer’s formula. The molar heat capacities of a gas at constant pressure
and constant volume are 28.8 J mol−1 K−1 and 20.5 J mol−1 K−1 respectively.
Calculate the gas constant.
1
Expert's answer
2018-03-24T07:51:07-0400
The Mayer’s law states
C_(P,m)-C_(V,m)=R
Therefore, the gas constant
R=28.8-20.5=8.3 J/(mol∙K)